OpenStudy (rainbow_dash):

What is the equation of a circle with center (−3, 0) and radius 20?

OpenStudy (anonymous):

the general equation of circle with center a,b) and radius r is \[\Large (x-a)^2+(y-b)^2=r^2\] in your case use a=-3 ,b=0 and r=20 can you find it now?

OpenStudy (anonymous):

@Rainbow_Dash waiting for your reply !!! did you get it ?

OpenStudy (rainbow_dash):

oookay yeah, thank you!